OpenStudy (anonymous):

Evaluate Arccsc 2. A. Pi/3 B. { pi/6, 5pi/6} C. { pi/6 +/- 2pin, 5pi/6 +/- 2pin} D. {pi/3 +/- 2pin, 5pi/3 +/- 2pin} ***My Answer: C*** Is that correct?

OpenStudy (anonymous):

arccsc 2 has the same meaning as 1/ sin x = 2 sinx =1/2 just keep in mind arc csc and arc sin they have same range so we will not lose any answer if sinx =1/2 then arc sin (1/2) =? arc sin (1/2) = 30 degree or Pi/6

OpenStudy (anonymous):

Ok so if arc sin (1/2) = pi/6, that means we can take out A and D right?

OpenStudy (anonymous):

The answer is C

OpenStudy (anonymous):

yes ...you are right